# A Process Mining-Based System For The Analysis and Prediction of Software Development Workflows

### v1

@@ 1 Introduction
In modern software engineering, DevOps practices have become the cornerstone of achieving continuous integration and continuous delivery (CI/CD). However, as software systems and teams scale, the complexity of development pipelines increases, making it difficult to maintain visibility over the entire workflow, from initial code contributions to deployment in production environments. While DevOps metrics, such as those popularized by the DORA framework [KF20], provide a valuable high-level overview of performance, they fail to capture the fine-grained process dynamics that lead to inefficiencies or bottlenecks in practice.
@@ 2 Related Work
DevOps research has been largely focused on measuring performance and organizational outcomes through key indicators such as deployment frequency, lead time for changes, change failure rate, and time to restore service [FHK18]. These so-called DORA metrics have become the industry standard for assessing software delivery performance [KF20]. However, while they provide high-level benchmarks, several authors [ESS23, KES22] have highlighted their limitations in explaining the underlying causes of inefficiencies or delays. They are primarily outcome-based and lack the granularity required to understand the behavioral and procedural factors that shape these outcomes.
Process Mining [vdA16] has proven to be an effective method for discovering, monitoring, and improving processes using event data extracted from information systems. Its application in the domain of software engineering is still emerging but promising. Recent studies [MCRRC21, CDLRR22] have shown that process mining can uncover complex workflows within software development environments, revealing handovers, rework loops, and deviations from expected paths. Despite this potential, practical applications in CI/CD pipelines remain limited due to the lack of standardized event data structures and the fragmented nature of development toolchains.
In recent years, research on Predictive Process Monitoring (PPM) has explored the use of machine learning techniques - such as gradient boosting, random forests, and deep neural networks - to predict process outcomes such as remaining time, success probability or compliance violations [TVLRD17, DFDMT19]. These techniques have demonstrated the ability to anticipate process behavior by learning from historical execution data. When applied to DevOps environments, similar approaches can be leveraged to forecast delivery delays or detect process anomalies [KBD23]. Nevertheless, few works have integrated predictive analytics with real, event-level DevOps data in a production context.

### v2

@@ 1 Introduction
In modern software engineering, DevOps practices have become the cornerstone of achieving continuous integration and continuous delivery (CI/CD). However, as software systems and teams scale, the complexity of development pipelines increases, making it difficult to maintain visibility over the entire workflow, from initial code contributions to deployment in production environments. While DevOps metrics, such as those popularized by the DORA framework [DR23], provide a valuable high-level overview of performance, they fail to capture the fine-grained process dynamics that lead to inefficiencies or bottlenecks in practice.
@@ 2 Related Work
DevOps research has been largely focused on measuring performance and organizational outcomes through key indicators such as deployment frequency, lead time for changes, change failure rate, and time to restore service [FHK18]. These so-called DORA metrics have become the industry standard for assessing software delivery performance. However, while providing high-level benchmarks, these can be misused [DR23] and are limited in explaining the underlying causes of inefficiencies or delays. They are primarily outcome-based and lack the granularity required to understand the behavioral and procedural factors that shape these outcomes.
Process Mining [vdA16] has proven to be an effective method for discovering, monitoring, and improving processes using event data extracted from information systems. Its application in the domain of software engineering is still emerging but promising. Recent studies [NZR24] have shown that process mining can uncover complex workflows within software development environments, revealing handovers, rework loops, and deviations from expected paths. Despite this potential, practical applications in CI/CD pipelines remain limited due to the lack of standardized event data structures and the fragmented nature of development toolchains.
In recent years, research on Predictive Process Monitoring (PPM) has explored the use of machine learning techniques - such as gradient boosting, random forests, and deep neural networks - to predict process outcomes such as remaining time, success probability or compliance violations [TVLRD17, CCDW+24]. These techniques have demonstrated the ability to anticipate process behavior by learning from historical execution data. When applied to DevOps environments, similar approaches can be leveraged to forecast delivery delays or detect process anomalies [BP24]. Nevertheless, few works have integrated predictive analytics with real, event-level DevOps data in a production context.